# Integrable measure equivalence rigidity of right-angled Artin groups via quasi-isometry

### v1

@@ 10 Lattice embeddings of right-angled Artin groups
In this section, we first prove a theorem that gives restrictions on the possible lattice embeddings of a countable group $H$ with bounded torsion which is measure equivalent to a right-angled Artin group $G$ with $|\operatorname{Out}(G)|<+\infty$ (Theorem 10.1). And we then describe all possible lattice embeddings under an integrability condition on the coupling between $G$ and $H$ (Theorem 10.5). For the second statement, we will first need to introduce the language of quasi-actions, which will be also useful in the next section.
@@ 10.1 Lattice embeddings and measure equivalence
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $H$ be a countable group with bounded torsion that is measure equivalent to $G$. Let $\mathfrak{H}$ be a locally compact second countable group, and let $\tau:H\to\mathfrak{H}$ be a lattice embedding.
@@
The intersection $H\cap K$ is finite, whence a lattice in the compact group $K$. Let $\pi:\mathfrak{H}\to\mathfrak{H}/K$ be the projection map. By [Rag72, Theorem 1.13] (as restated in [BFS20, Lemma 3.5]), the $\pi$-image of $H$ in $\mathfrak{H}/K$ is again a lattice. Since $\mathrm{Aut}(\mathbb{B})$ is totally disconnected, so is its closed subgroup $\mathfrak{H}/K$. Since every lattice embedding of a countable group with bounded torsion in a totally disconnected locally compact group is cocompact (see [BCGM19, Corollary 4.11]), it follows that the image of $H$ in $\mathfrak{H}/K$ is cocompact. Let $K^{\prime}\subseteq\mathfrak{H}/K$ be a compact set whose $H$-translates cover $\mathfrak{H}/K$. The map $\pi$ is closed, continuous, and it has compact fibers because $K$ is compact, see [AT08, Theorem 1.5.7]. It follows that $\pi^{-1}(K^{\prime})$ is compact, see [Eng89, Theorem 3.7.2], and therefore $H$ is cocompact in $\mathfrak{H}$. ∎
@@ 10.2 Quasi-actions
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$. Let $H$ be a finitely generated group that is quasi-isometric to $G$, and let $\mathfrak{H}$ be a locally compact second countable group that contains $H$ as a cocompact lattice.
@@
By [MSW03, Lemma 28], the group $\mathfrak{H}$ has a proper, cobounded, measurable quasi-action $\rho$ on $H$. Notice that, while properness and measurability of the quasi-action are not stated in [MSW03, Lemma 28], they follow from the construction. Indeed, to construct $\rho$, one starts with an enumeration $H=\{h_{n}\}_{n\in\mathbb{N}}$ and chooses a symmetric compact subset $K\subseteq\mathfrak{H}$ containing the identity, such that $HK=\mathfrak{H}$. For $g\in\mathfrak{H}$ and $h\in H$, one lets $\rho(g,h)=h^{\prime}$, where $h^{\prime}$ is the first element of the enumeration of $H$ such that $gh\in h^{\prime}K$. In particular $\mathfrak{H}$ has a proper, cobounded, measurable quasi-action on $G$, through a quasi-isometry between $H$ and $G$.
By Corollary 10.3, there exists a unique continuous flat-preserving action $\alpha:\mathfrak{H}\to\mathrm{Bij}_{\mathrm{FP}}(G)$ such that there exists $D\geq 0$ such that for every $h\in\mathfrak{H}$ and every $g\in G$, one has $d(\rho(h,g),\alpha(h)(g))\leq D$. In particular the action $\alpha$ is by quasi-isometries, and it is proper and cobounded. This proves the first part of the proposition.
The second part of the proposition then follows from [HK18, Theorem 6.2], and $Y$ arises as a blow-up building as in Section 3. As in the proof of [HK18, Theorem 6.2], the properness of the $\mathfrak{H}$-action on $Y$ follows from the properness of its action on $G$, and the fact that there is an equivariant quasi-isometry from $G$ to the set of rank 00 vertices of $Y$ (see the paragraphs after the proof of the claim in [HK18, p. 587]). The coboundedness (in fact cocompactness as $Y$ is locally finite) of the action follows from the same argument. And the measurability of the $\mathfrak{H}$-action on $Y$ also follows from the measurability of its action on $G$ by the same argument (the argument gives the measurability at the level of rank 00 vertices, but the action on rank 00 vertices determines the whole action). Finally, the continuity of the homomorphism $\mathfrak{H}\to\mathrm{Aut}(Y)$ follows again from an automatic continuity statement due to Mackey, see [Zim84, Theorem B.3]. ∎
@@ 10.3 Lattice embeddings under an integrability assumption
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $H$ be a countable group with bounded torsion. Assume that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$. Let $\mathfrak{H}$ be a locally compact second countable group in which $H$ embeds as a lattice.
@@ 11.1 Two arguments for extending actions to a bigger complex
For each $n\in\mathbb{N}$, let $T_{n}$ be the regular tree of valence $n$, with each edge of length $2$. An action of a group $H$ on $T_{n}$ by automorphisms is even if it preserves the colours of vertices in any bipartite colouring of $VT_{n}$.
@@
Let $S=\{s_{1},\dots,s_{n}\}$ be a finite set of cardinality $n$. We label each edge of $T_{n}$ by a letter in $S$, in such a way that two edges incident on the same vertex never have the same label. We fix a bipartition of $T_{n}$ into black and white vertices. Let $T^{\prime}_{n}$ be the barycentric subdivision of $T_{n}$. Vertices of $T^{\prime}_{n}$ are of three types: black vertices (in $V(T_{n})$), white vertices (in $V(T_{n})$), and gray vertices, corresponding to the midpoints of edges of $T_{n}$. We orient each edge of $T^{\prime}_{n}$ so that it points towards its gray vertex. We fix a set $\{a_{1},\dots,a_{n},a^{\prime}_{1},\dots,a^{\prime}_{n}\}$ of labels. Any edge of $T^{\prime}_{n}$ from a white (resp. black) vertex of $T_{n}$ to the midpoint of an $s_{i}$-labeled edge of $T_{n}$, is labeled by $a_{i}$ (resp. $a^{\prime}_{i}$). Let $W_{2n}$ be a wedge sum of $2n$ oriented circles, labeled by $a_{1},\ldots,a_{n},a^{\prime}_{1},\ldots,a^{\prime}_{n}$. There is a unique map $\pi:T^{\prime}_{n}\to W_{2n}$ which preserves labels and orientations of edges. But $\pi$ is not a covering map. Now we enlarge $T^{\prime}_{n}$ to a larger space $K_{n}$, and extend $\pi$ to a map $\pi:K_{n}\to W_{2n}$ which is a covering map. This relies on Haglund and Wise’s canonical completion procedure [HW08, Section 6], which we explain in this special case. We enlarge $T^{\prime}_{n}$ as follows. For each edge $e$ of $T^{\prime}_{n}$ oriented from a vertex $x$ to another vertex $y$, we add an edge $e^{\prime}$ to $T^{\prime}_{n}$ such that $e^{\prime}$ and $e$ have the same label, but $e^{\prime}$ is oriented from $y$ to $x$. Next for each white (resp. black) vertex $x$ of $T^{\prime}_{n}$, we attach $n$ oriented loops based $x$, labeled by $\{a^{\prime}_{1},\ldots,a^{\prime}_{n}\}$ (resp. $\{a_{1},\ldots,a_{n}\}$). Finally for each gray vertex $x$ of $T^{\prime}_{n}$ which is the midpoint of an $s_{i}$-labeled edge of $T_{n}$, we attached $2(n-1)$ oriented loops based at $x$, labeled by $\{a_{1},\ldots,\hat{a}_{i},\ldots,a_{n},a^{\prime}_{1},\ldots,\hat{a}^{\prime}_{i},\ldots,a^{\prime}_{n}\}$ – here $\hat{a}_{i},\hat{a}^{\prime}_{i}$ means that we remove these two elements from the set of labels. Now one readily verifies that the map $\pi:T^{\prime}_{n}\to W_{2n}$ extends to a label and orientation preserving covering map $\pi:K_{n}\to W_{2n}$. A standard circle in $K_{n}$ is an embedded copy of $\mathbb{S}^{1}$ made of edges with the same label. Each standard circle in $K_{n}$ has either one or two edges. Now consider the action $H\curvearrowright T_{n}\times T_{n}$, which gives two factor actions $H\curvearrowright T_{n}$ preserving the bipartite vertex coloring of $T_{n}$. Any factor action induces an action $H\curvearrowright T^{\prime}_{n}$ which preserves orientation of edges and colors of vertices of $T^{\prime}_{n}$ (though it might not preserve labels of edges of $T^{\prime}_{n}$). And since the number of edge-loops at vertices of a given colour is constant, the action $H\curvearrowright T^{\prime}_{n}$ extends to an action $H\curvearrowright K_{n}$ which preserves the orientation of edges and sends standard circles to standard circles. Now the free and cocompact action $H\curvearrowright T_{n}\times T_{n}$ extends to a free and cocompact action $H\curvearrowright K_{n}\times K_{n}$. A standard torus in $K_{n}\times K_{n}$ is a product of two standard circles, one from each factor. Then the action $H\curvearrowright K_{n}\times K_{n}$ preserves the edge orientation and sends every standard torus to a standard torus. Moreover, the subspace $T^{\prime}_{n}\times T^{\prime}_{n}$ is invariant under this action. Let $J_{2n}\times J_{2n}$ be the universal cover of $K_{n}\times K_{n}$ with the induced label and orientation of edges (where the edges coming from the second factor are labeled by $\tilde{a}_{i},\tilde{a}^{\prime}_{i}$). Then standard flats in $J_{2n}\times J_{2n}$ are lifts of standard tori in $K_{n}\times K_{n}$ – for this it is important to notice that components of $K_{n}$ consisting of edges with the same label are reduced to circles. Let $H^{\prime}\subseteq\mathrm{Aut}(J_{2n}\times J_{2n})$ be the subgroup consisting of all automorphisms that lift automorphisms in $H$. Then the $H^{\prime}$-action on $J_{2n}\times J_{2n}$ is free and cocompact, preserves the orientation of edges, and sends standard flats to standard flats. In addition, the inclusion map $T^{\prime}_{n}\times T^{\prime}_{n}\to K_{n}\times K_{n}$ lifts to an isometric embedding $\theta:T^{\prime}_{n}\times T^{\prime}_{n}\to J_{2n}\times J_{2n}$. As $T^{\prime}_{n}\times T^{\prime}_{n}$ is invariant under the $H$-action on $K_{n}\times K_{n}$, such a lift gives an injective homomorphism $\varphi:H\to H^{\prime}$ such that $\theta$ is equivariant with respect to $\varphi$. Finally, the additional part of the lemma follows from our construction. ∎
@@
We need another extension criterion, which is based on an idea of Hughes in [Hug21, Section 7.2]. Our next lemma is a variation over a statement to be found in ongoing work of Mj and the second-named author [HM].
@@
Let $\Gamma_{1}$ be an induced subgraph of $\Gamma_{2}$. Suppose $H$ is a group acting freely and cocompactly on $X_{\Gamma_{1}}$, preserving the orientation of edges, and sending standard flats to standard flats. Let $c_{1}:H\times VX_{\Gamma_{1}}\to\mathrm{Aut}(\Gamma_{1})$ be the associated type cocycle. Suppose that there exists a cocycle $c_{2}:H\times VX_{\Gamma_{1}}\to\mathrm{Aut}(\Gamma_{2})$ such that $c_{2}(h,x)_{\mid\Gamma_{1}}=c_{1}(h,x)$ for any $h\in H$ and $x\in VX_{\Gamma_{1}}$. Then there exist a group $H^{\prime}$ acting freely and compactly on $X_{\Gamma_{2}}$ sending standard flats to standard flats, an injective group homomorphism $\phi:H\to H^{\prime}$, and a $\phi$-equivariant embedding $j:X_{\Gamma_{1}}\to X_{\Gamma_{2}}$ preserving labels and orientations of edges. Moreover, each element of $H^{\prime}$ sends standard lines labeled by $v\in V\Gamma_{2}$ to standard lines whose labels belong to the orbit of $v$ under the action of elements in $\{c_{2}(h,x)\}_{h\in H,x\in VX_{\Gamma_{1}}}$.
@@ 11.2 On local actions
The following can be deduced from work of Lazarovich–Levcovitz–Margolis [LLM22], relying on earlier works by Burger–Mozes [BM00a, BM00b] and Radu [Rad20].
@@
By [LLM22, Theorem 4.2 and Lemma 4.3], there is a group $H$ acting on $T_{n}\times T_{n}$ as a uniform lattice such that all the local groups are the full symmetry groups on $n$ letters. Moreover, $H$ has an index $4$ subgroup $H^{+}$ which is simple and preserves the bipartition of each factor tree (see the beginning of [LLM22, Section 2] for the definition of $H^{+}$). In addition $H^{+}$ is torsion-free by [Rad20, Lemma 3.1]. Thus the group $H^{+}$ satisfies the requirement of the lemma. ∎

### v2

@@ 10 Lattice embeddings of right-angled Artin groups
In this section, we first prove a theorem that gives restrictions on the possible lattice embeddings of a countable group $H$ with bounded finite subgroups which is measure equivalent to a right-angled Artin group $G$ with $|\operatorname{Out}(G)|<+\infty$ (Theorem 10.1). And we then describe all possible lattice embeddings under an integrability condition on the coupling between $G$ and $H$ (Theorem 10.5). For the second statement, we will first need to introduce the language of quasi-actions, which will be also useful in the next section.
@@ 10.1 Lattice embeddings and measure equivalence
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $H$ be a countable group with bounded finite subgroups that is measure equivalent to $G$. Let $\mathfrak{H}$ be a locally compact second countable group, and let $\tau:H\to\mathfrak{H}$ be a lattice embedding.
@@
The intersection $H\cap K$ is finite, whence a lattice in the compact group $K$. Let $\pi:\mathfrak{H}\to\mathfrak{H}/K$ be the projection map. By [Rag72, Theorem 1.13] (as restated in [BFS20, Lemma 3.5]), the $\pi$-image of $H$ in $\mathfrak{H}/K$ is again a lattice. Since $\mathrm{Aut}(\mathbb{B})$ is totally disconnected, so is its closed subgroup $\mathfrak{H}/K$. Since every lattice embedding of a countable group with bounded finite subgroups in a totally disconnected locally compact group is cocompact (see [BCGM19, Corollary 4.11]), it follows that the image of $H$ in $\mathfrak{H}/K$ is cocompact. Let $K^{\prime}\subseteq\mathfrak{H}/K$ be a compact set whose $H$-translates cover $\mathfrak{H}/K$. The map $\pi$ is closed, continuous, and it has compact fibers because $K$ is compact, see [AT08, Theorem 1.5.7]. It follows that $\pi^{-1}(K^{\prime})$ is compact, see [Eng89, Theorem 3.7.2], and therefore $H$ is cocompact in $\mathfrak{H}$. ∎
@@ 10.2 Quasi-actions
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$. Let $H$ be a finitely generated group that is quasi-isometric to $G$, and let $\mathfrak{H}$ be a compactly generated locally compact second countable group which is quasi-isometric to $G$.
@@
Let $X$ be a Cayley-Abels graph of $\mathfrak{H}$, equipped with the path metric with edge length $1$, and let $\varphi:X\to G$ be a quasi-isometry, with its quasi-inverse given by $\psi:G\to X$. Then the proper and cocompact isometric action $\rho:\mathfrak{H}\times X\to X$ gives a proper and cobounded quasi-action $\rho^{\prime}:\mathfrak{H}\times G\to G$ with $\rho^{\prime}(\cdot,g)=\varphi(\rho(\cdot,\psi(g)))$.
By Corollary 10.3, there exists a unique continuous flat-preserving action $\alpha:\mathfrak{H}\to\mathrm{Bij}_{\mathrm{FP}}(G)$ such that there exists $D\geq 0$ such that for every $h\in\mathfrak{H}$ and every $g\in G$, one has $d(\rho^{\prime}(h,g),\alpha(h)(g))\leq D$. In particular the action $\alpha$ is by quasi-isometries, and it is proper and cobounded. This proves the first part of the proposition.
The second part of the proposition then follows from [HK18, Theorem 6.2], and $Y$ arises as a blow-up building as in Section 3. As in the proof of [HK18, Theorem 6.2], the properness of the $\mathfrak{H}$-action on $Y$ follows from the properness of its action on $G$, and the fact that there is an equivariant quasi-isometry from $G$ to the set of rank 0 vertices of $Y$ (see the paragraphs after the proof of the claim in [HK18, p. 587]). The coboundedness (in fact cocompactness as $Y$ is locally finite) of the action follows from the same argument. The continuity of the action follows from [Cor18, Lemma 19.29]. ∎
@@ 10.3 Lattice embeddings under an integrability assumption
Let $G$ be a non-cyclic right-angled Artin group with $|\operatorname{Out}(G)|<+\infty$, and let $H$ be a countable group with bounded finite subgroups. Assume that there exists an $(L^{1},L^{0})$-measure equivalence coupling from $H$ to $G$. Let $\mathfrak{H}$ be a locally compact second countable group in which $H$ embeds as a lattice.
@@ 11.1 Two arguments for extending actions to a bigger complex
We make the following definition.
Definition 11.1 (Even action on a tree).
An action of a group $H$ on a tree $T$ is even if it preserves the colours of vertices in any bipartite colouring of the vertex set $VT$.
For each $n\in\mathbb{N}$, let $T_{n}$ be the regular tree of valence $n$, with each edge of length $2$.
@@
Let $S=\{s_{1},\dots,s_{n}\}$ be a finite set of cardinality $n$. We label each edge of $T_{n}$ by a letter in $S$, in such a way that two edges incident on the same vertex never have the same label. We fix a bipartition of $T_{n}$ into black and white vertices. Let $T^{\prime}_{n}$ be the barycentric subdivision of $T_{n}$. Vertices of $T^{\prime}_{n}$ are of three types: black vertices (in $V(T_{n})$), white vertices (in $V(T_{n})$), and gray vertices, corresponding to the midpoints of edges of $T_{n}$. We orient each edge of $T^{\prime}_{n}$ so that it points towards its gray vertex. We fix a set $\{a_{1},\dots,a_{n},a^{\prime}_{1},\dots,a^{\prime}_{n}\}$ of labels. Any edge of $T^{\prime}_{n}$ from a white (resp. black) vertex of $T_{n}$ to the midpoint of an $s_{i}$-labeled edge of $T_{n}$, is labeled by $a_{i}$ (resp. $a^{\prime}_{i}$).
Let $W_{2n}$ be a wedge sum of $2n$ oriented circles, labeled by $a_{1},\ldots,a_{n},a^{\prime}_{1},\ldots,a^{\prime}_{n}$. There is a unique map $\pi:T^{\prime}_{n}\to W_{2n}$ which preserves labels and orientations of edges. But $\pi$ is not a covering map. Now we enlarge $T^{\prime}_{n}$ to a larger space $K_{n}$, and extend $\pi$ to a map $\pi:K_{n}\to W_{2n}$ which is a covering map. This relies on Haglund and Wise’s canonical completion procedure [HW08, Section 6], which we explain in this special case.
We enlarge $T^{\prime}_{n}$ as follows. For each edge $e$ of $T^{\prime}_{n}$ oriented from a vertex $x$ to another vertex $y$, we add an edge $e^{\prime}$ to $T^{\prime}_{n}$ such that $e^{\prime}$ and $e$ have the same label, but $e^{\prime}$ is oriented from $y$ to $x$. Next for each white (resp. black) vertex $x$ of $T^{\prime}_{n}$, we attach $n$ oriented loops based $x$, labeled by $\{a^{\prime}_{1},\ldots,a^{\prime}_{n}\}$ (resp. $\{a_{1},\ldots,a_{n}\}$). Finally for each gray vertex $x$ of $T^{\prime}_{n}$ which is the midpoint of an $s_{i}$-labeled edge of $T_{n}$, we attached 2(n1)2(n-1) oriented loops based at $x$, labeled by $\{a_{1},\ldots,\hat{a}_{i},\ldots,a_{n},a^{\prime}_{1},\ldots,\hat{a}^{\prime}_{i},\ldots,a^{\prime}_{n}\}$ – here $\hat{a}_{i},\hat{a}^{\prime}_{i}$ means that we remove these two elements from the set of labels. Now one readily verifies that the map $\pi:T^{\prime}_{n}\to W_{2n}$ extends to a label and orientation preserving covering map $\pi:K_{n}\to W_{2n}$.
A standard circle in $K_{n}$ is an embedded copy of $\mathbb{S}^{1}$ made of edges with the same label. Each standard circle in $K_{n}$ has either one or two edges.
Now consider the action $H\curvearrowright T_{n}\times T_{n}$, which gives two factor actions $H\curvearrowright T_{n}$ preserving the bipartite vertex coloring of $T_{n}$. Any factor action induces an action $H\curvearrowright T^{\prime}_{n}$ which preserves orientation of edges and colors of vertices of $T^{\prime}_{n}$ (though it might not preserve labels of edges of $T^{\prime}_{n}$). And since the number of edge-loops at vertices of a given colour is constant, the action $H\curvearrowright T^{\prime}_{n}$ extends to an action $H\curvearrowright K_{n}$ which preserves the orientation of edges and sends standard circles to standard circles.
Now the free and cocompact action $H\curvearrowright T_{n}\times T_{n}$ extends to a free and cocompact action $H\curvearrowright K_{n}\times K_{n}$. A standard torus in $K_{n}\times K_{n}$ is a product of two standard circles, one from each factor. Then the action $H\curvearrowright K_{n}\times K_{n}$ preserves the edge orientation and sends every standard torus to a standard torus. Moreover, the subspace $T^{\prime}_{n}\times T^{\prime}_{n}$ is invariant under this action. Let $J_{2n}\times J_{2n}$ be the universal cover of $K_{n}\times K_{n}$ with the induced label and orientation of edges (where the edges coming from the second factor are labeled by $\tilde{a}_{i},\tilde{a}^{\prime}_{i}$). Then standard flats in $J_{2n}\times J_{2n}$ are lifts of standard tori in $K_{n}\times K_{n}$ – for this it is important to notice that components of $K_{n}$ consisting of edges with the same label are reduced to circles.
Let $H^{\prime}\subseteq\mathrm{Aut}(J_{2n}\times J_{2n})$ be the subgroup consisting of all automorphisms that lift automorphisms in $H$. Then the $H^{\prime}$-action on $J_{2n}\times J_{2n}$ is free and cocompact, preserves the orientation of edges, and sends standard flats to standard flats. In addition, the inclusion map $T^{\prime}_{n}\times T^{\prime}_{n}\to K_{n}\times K_{n}$ lifts to an isometric embedding $\theta:T^{\prime}_{n}\times T^{\prime}_{n}\to J_{2n}\times J_{2n}$. As $T^{\prime}_{n}\times T^{\prime}_{n}$ is invariant under the $H$-action on $K_{n}\times K_{n}$, such a lift gives an injective homomorphism $\varphi:H\to H^{\prime}$ such that $\theta$ is equivariant with respect to $\varphi$. Finally, the additional part of the lemma follows from our construction. ∎
@@
We need another extension criterion, which is based on an idea of Hughes in [Hug21, Section 7.2]. Our next lemma is a variation over a statement from work of Mj and the second-named author [HM23, Section 4.2].
@@
Let $\Gamma_{1}$ be an induced subgraph of $\Gamma_{2}$. Suppose $H$ is a group acting freely and cocompactly on $X_{\Gamma_{1}}$, preserving the orientation of edges, and sending standard flats to standard flats. Let $c_{1}:H\times VX_{\Gamma_{1}}\to\mathrm{Aut}(\Gamma_{1})$ be the associated type cocycle. Suppose that there exists a cocycle $c_{2}:H\times VX_{\Gamma_{1}}\to\mathrm{Aut}(\Gamma_{2})$ such that $c_{2}(h,x)_{\mid\Gamma_{1}}=c_{1}(h,x)$ for any $h\in H$ and $x\in VX_{\Gamma_{1}}$.
Then there exist a group $H^{\prime}$ acting freely and compactly on $X_{\Gamma_{2}}$ sending standard flats to standard flats, an injective group homomorphism $\phi:H\to H^{\prime}$, and a $\phi$-equivariant embedding $j:X_{\Gamma_{1}}\to X_{\Gamma_{2}}$ preserving labels and orientations of edges. Moreover, each element of $H^{\prime}$ sends standard lines labeled by $v\in V\Gamma_{2}$ to standard lines whose labels belong to the orbit of $v$ under the action of elements in $\{c_{2}(h,x)\}_{h\in H,x\in VX_{\Gamma_{1}}}$.
@@ 11.2 On local actions
The following can be deduced from work of Lazarovich–Levcovitz–Margolis [LLM23], relying on earlier works by Burger–Mozes [BM00a, BM00b] and Radu [Rad20] (recall Definition 11.1 for the notion of even factor actions).
@@
By [LLM23, Theorem 4.2 and Lemma 4.3], there is a group $H$ acting on $T_{n}\times T_{n}$ as a uniform lattice such that all the local groups are the full symmetry groups on $n$ letters. Moreover, $H$ has an index $4$ subgroup $H^{+}$ which is simple and preserves the bipartition of each factor tree (see the beginning of [LLM23, Section 2] for the definition of $H^{+}$). In addition $H^{+}$ is torsion-free by [Rad20, Lemma 3.1]. Thus the group $H^{+}$ satisfies the requirement of the lemma. ∎